OpenStudy (anonymous):

how do i solve the following set of equations? -x+2y+z+w=0 x+y-z+w=0 -x+y+z=0

hero (hero):

There are several ways to do it. Usually, they make you do it the hard way first, then show you the easy ways later.

hero (hero):

1. Elimination 2. Substitution 3. Matrices (Gaussian Elimination) 4. Cramer's Rule

OpenStudy (zzr0ck3r):

what class is this for?

hero (hero):

Oh wait, I just noticed all of them equal zero

OpenStudy (anonymous):

i tried using matrices but it seems impossible to get the variable values since i have 3 equations but 4 unknowns

hero (hero):

For that, there's this: http://www.gregthatcher.com/Mathematics/GaussJordan.aspx

OpenStudy (zzr0ck3r):

you will have a free variable?

OpenStudy (anonymous):

what is free variable?

OpenStudy (zzr0ck3r):

x = z y=0 w=0

hero (hero):

A free variable is basically what you get since w = 0 as implied by the third row

OpenStudy (zzr0ck3r):

is this for linear algebra?

OpenStudy (zzr0ck3r):

so i row reduced and found x = w ,y = 0 and z = 0, but im not sure if this is what you guys are doing in class.

OpenStudy (anonymous):

@zzr0ck3r ,yes it is linear algebra

OpenStudy (zzr0ck3r):

so when you row reduce you get 1 0 -1 0 0 0 1 0 0 0 0 0 0 1 0 right?

OpenStudy (anonymous):

from my augmented matrix i get w=0 y=0 x-y=0 which means x=0

OpenStudy (anonymous):

@zzr0ck3r, yes i get that

OpenStudy (zzr0ck3r):

so first row x y z w 1 0 -1 0 0 so x 0 -z 0 0 so x = z

OpenStudy (anonymous):

actually it's x-z=0 not x-y=0

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

so x = z

OpenStudy (zzr0ck3r):

so like you said x = z, y = 0, w = 0

OpenStudy (zzr0ck3r):

in other words, there are infinite many solutions.

OpenStudy (anonymous):

thanks a lot.@Hero and @zzr0ck3r

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

thanks @Hero